OpenStudy (anonymous):

Carissa and Louann both have savings accounts. Carissa has $250 in her account and deposits $80 per month. Louann has $1230 in her account and withdraws $60 per month. After how many months will they have equal amounts of money in their accounts? What amount of money will each have when it is the same amount? It will take months. Then, they will both have $ in their accounts.

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@DanielSherron

OpenStudy (tkhunny):

Month 0 250 1230 Month 1 250+80 = 330 1230-60 = 1170 Month 2 330+80 = 410 1170-60 = 1110 Are we supposed to be doing something harder?